# UniPart: Part-Level 3D Generation with Unified 3D Geom–Seg Latents

### v1

@@ 1 Introduction
With the rapid development of 3D content creation in fields such as computer graphics, robotics, virtual reality, and digital manufacturing, the demand for controllable and semantically structured 3D shape generation has increased significantly. Traditional 3D generation methods [zhang2024clay, zhao2025hunyuan3d2, li2025triposg, xiang2024trellis] primarily focus on producing holistic meshes that capture only the overall geometry of objects. However, real-world applications—such as part-aware editing, physical simulation, robotic manipulation, and modular design—require not only accurate geometry but also a decomposable understanding of object parts and their relationships. This need has motivated increasing interest in part-level 3D shape generation, which aims to synthesize objects as sets of coherent and semantically meaningful components. Compared to holistic 3D reconstruction, part-level 3D shape generation poses two principal challenges. First, achieving structurally and semantically coherent part segmentation is inherently challenging. The model must infer an appropriate segmentation granularity while preserving both global structural integrity and semantic consistency. The wide variability in object geometries and part compositions across—and even within—categories further makes coherent and meaningful decomposition highly nontrivial. Second, part geometry degradation frequently occurs during generation, particularly for small or thin components. These fine structures tend to lose geometric fidelity due to limited generative resolution and the difficulty of jointly preserving global coherence and local geometric details.
@@
Existing solutions develop implicit segmentation by composing groups of latent vectors for decomposable part generation [lin2025partcrafter, tang2025partpacker, chen2025autopartgen], but suffer from limited controllability over segmentation granularity and suboptimal part-level geometric fidelity. Another line of work first performs whole-object generation and subsequently applies explicit segmentation using multi-view images [chen2025partgen], dense points [yan2025xpart], low-resolution voxels [yang2025omnipart], or mesh faces [yang2025holopart]. These approaches achieve higher segmentation quality and controllability but require a powerful auxiliary segmenter trained on abundant labeled data. We begin with the observation that part awareness emerges implicitly during whole-geometry learning. This arises from two factors: (i) intra- and inter-part geometric correlations are implicitly promoted during VAE latent-space construction for efficient geometric compression, leading to part-aware latent representations; (ii) image conditions encoded by pretrained vision models such as DINOv2 [oquab2024dinov2] carry rich part-level semantic cues, which facilitate part-aware learning in image-to-geometry diffusion. Motivated by these observations, we propose to construct a unified 3D geometry–segmentation latent space based on the VecSet representation [zhang2024clay]. This latent space is learned to jointly encode geometric shape and part-segmentation information into Geom-Seg VecSet, and to decode each latent vector into its corresponding geometric contribution and part label. Thanks to the tight integration of part understanding and geometry learning, our representation maintains reconstruction quality comparable to the original VecSet.
@@
For image-guided part-level 3D generation, we introduce a two-stage latent diffusion model, UniPart. We first perform whole-object diffusion to generate Geom-Seg VecSet, yielding both the global geometric latent and the corresponding part latent mask. Compared with methods that conduct part-level generation with implicit segmentation, our approach produces explicit part-latent masks, providing more flexible segmentation granularity and higher part-level controllability. Unlike previous generate-then-segment pipelines, our approach performs geometry generation and part segmentation jointly. This enables more efficient part learning by leveraging the rich part-aware priors implicitly captured in the geometry generator. Moreover, latent-space segmentation actually strikes a balance between granularity—akin to dense point-based approaches—and robustness—similar to low-resolution voxel methods. In the second stage, we directly use the segmented part latents, along with the global latent, as dual conditioning signals for part-level generation. Compared with prior approaches that require whole-object geometry generation and segmentation on explicit representations to condition the part-level diffusion [yan2025xpart, yang2025holopart, yang2025omnipart], our latent-based conditioning offers three key advantages: (i) avoiding geometry degradation introduced by decode–encode cycles (e.g., decoding to SDF for lossy marching-cube mesh extraction and re-encoding sampled surface points [yan2025xpart, yang2025holopart]); (ii) providing domain-aligned conditioning signals that facilitate latent-space diffusion learning; (iii) injecting informative part–whole relational cues inherently preserved in the latent space. Furthermore, our part-level diffusion performs a novel dual-space latent generation procedure that jointly predicts part geometry latents in both the global coordinate space and a normalized canonical space. Decoding these two latents provides the part’s canonical geometry (from the normalized space latent) and its scale and placement within the object (from the global-space latent), which substantially mitigates part-geometry degradation and enables seamless and consistent part composition. Extensive experiments are conducted to validate the effectiveness and superiority of our UniPart on both part segmentation and geometry quality.
@@ 2 Related Work
Recent object-level 3D generation methods fall into two main categories: 2D-lifting and native 3D approaches. DreamFusion [poole2022dreamfusion] pioneered the use of 2D diffusion priors via Score Distillation Sampling [alldieck2024sds, wang2023sjc] to optimize 3D representations, while follow-ups [long2023wonder3d, wang2024prolificdreamer, liusyncdreamer, shi2023mvdream] improve multi-view consistency by generating aligned image sets before reconstruction. In parallel, native 3D generative models are trained directly on large-scale datasets [objaverse, deitke2024objaversexl]. CLAY [zhang2024clay] and related works [zhao2025hunyuan3d2, li2025triposg, li2024craftsman] adopt a VecSet-based VAE–LDM framework [zhang20233dshape2vecset] to generate meshes and require high-quality watertight meshes for training. Trellis [xiang2024structured] introduces an explicit voxel-based latent space trained with rendering losses, offering stronger geometric inductive bias [li2025sparc3d, he2025sparseflex]. However, scaling such representations to high resolution remains a key challenge for fine-detail synthesis. Our work operates at the part level and builds upon the VecSet paradigm: we retain the VecSet-based latent architecture but extend the purely geometric latent space by incorporating geometry-aligned segmentation features, thereby constructing a novel unified geometry–segmentation latent space that enables structured part decomposition while preserving reconstruction fidelity.
@@
Recent part segmentation methods leverage either 2D priors or native 3D supervision. Approaches such as SaMesh [tang2024samesh] lift 2D SAM [kirillov2023sam, ravi2024sam2] predictions to 3D via multi-view post-processing, while SAMPart3D [yang2024sampart3d] and PartField [liu2025partfield] distill SAM priors into a 3D semantic feature field for clustering-based decomposition. In contrast, 3D native methods like P3-SAM [ma2025p3sam] and PartSAM [zhu2025partsam] are trained directly on large-scale part-annotated meshes and produce point-wise semantic features for mask prompting. GeoSAM2 [deng2025geosam2] adopts a hybrid strategy: it finetunes the pretrained SAM on 3D data to enable multi-view consistent segmentation, then back-projects the 2D segmentation onto the mesh surface to obtain a coherent 3D part decomposition. Unlike these strategies that conduct segmentation on explicit 2D/3D representations, our method performs decomposition on the latent representation. By integrating high-dimensional, geometry-aligned semantic features into the latent representation, our approach inherits the robustness and boundary-awareness of sampling while enabling end-to-end part-level generation without relying on hard segmentation masks. Meanwhile, solving the segmentation task as a diffusion-sampling task is much more reasonable than regression. Unlike regression, which forces a single deterministic output, treating part segmentation as a diffusion-based sampling task better captures the inherent ambiguity and multi-modality of valid part decompositions.
@@
Existing part-level 3D generation methods mainly fall into two groups based on the diffusion and autoregressive paradigms. Diffusion-based approaches can be further categorized by their generation strategy: one-shot-for-all methods such as PartCrafter [lin2025partcrafter], PartPacker [tang2025partpacker], OmniPart [yang2025omnipart], BANG [zhang2025bang], FullPart [ding2025fullpart] and X-Part [yan2025xpart], which jointly generate all parts in a single forward pass; one-shot-per-part models such as HoloPart [yang2025holopart] generates each part independently but conditioned on global context, and AutoPartGen [chen2025autopartgen] adopts a sequential generation scheme. In contrast, AR-based methods operate directly on mesh sequences: MeshCoder [dai2025meshcoderllmpoweredstructuredmesh] leverages an LLM to produce Blender scripts that are executed to reconstruct parts, and Mesh Silksong [song2025meshsilksongautoregressivemesh] employs a manifold-aware mesh representation for autoregressive part synthesis. Except for adopting joint geometry-segmentation diffusion at the first stage, our part-level diffusion is also different in: (i) using latent-based conditions rather than explicit representations, which avoids geometry degradation, provides domain-aligned conditioning signals, and also injects informative part–whole relational cues; (ii) performing a dual-space latent generation procedure that achieves full-resolution part geometry of high quality and enables seamless and consistent part composition.
@@ 3.1 Geometry-Segmentation VAE
[zhang2024clay] is a popular representation that encodes the geometry of a 3D mesh into the latent space. Specifically, dense points $P\in\mathbb{R}^{C\times 6}$ (each of $C$ points is represented by its position and normal vector on the mesh surface) are first sampled from the mesh $\mathcal{O}$. A fixed-length set of latent vectors $Z\in\mathbb{R}^{L\times d}$ is computed through a cross-attention layer between $P$ and its downsampled $L$ points (or a learnable query set) and several self-attention layers, encoded into $d$-dimension latent vectors. This process is conducted by the encoder $\mathcal{E}$ of a transformer-based VAE. The decoder $\mathcal{D}$ then decodes $Z$ into an implicit field (e.g. occupancy field or signed-distance field) for mesh reconstruction, as $\mathcal{D}(Z,q)\rightarrow[0,1]$, where $q$ is the set of query points in the target 3D space. The whole VAE is trained by minimizing the loss function formulated as
@@
Classic VecSet representations encode rich geometric information into the latent space. To better support compositional generation, we further propose Geom-Seg Vecset that additionally encodes part segmentation information into the latent space and conducts latent segmenting in the decoding stage for part-wise representation. In doing so, we first segment object meshes into $N$ parts according to both geometric and semantic patterns. Then $C$ points with part segmentation label are sampled, denoted as $P\in\mathbb{R}^{C\times 7}$, as the input of encoding, derieving the encoded latent $Z=\mathcal{E}(P)\in\mathbb{R}^{L\times d}$. A decoder $\mathcal{D}_{\text{geom}}$ as in VecSet is adopted for geometry reconstruction from $Z$. We further develop a segmentation decoder $\mathcal{D}_{\text{seg}}$ and follow the promptable segmentation pipeline of SAM2 [ravi2024sam2], which accepts $Z$ and latent indices $q$ as input and generates segmentation masks that the prompt belongs to. The Geom-Seg VecSet VAE is trained by minimizing the loss function formulated as
@@
where $L_{\text{seg}}$ is the segmentation loss following SAM2 [ravi2024sam2]. Considering geometric cues often delineate semantic boundaries, the geometric features and part segmentation patterns of objects are inherently interdependent. It critically facilitates coupled representation learning of geometry and segmentation information. In practice, we found that the Geom-Seg VecSet space can be smoothly learned by fine-tuning a pure geometric VecSet VAE, with the latent space and neural network kept in the original scale (equal in $L$ and $d$), which is sufficient to jointly represent the segmentation and geometry without loss of quality (see visualizations in Sec. 4.3).
@@ 3.2 Whole-level Latent Diffusion
With the learned geometry-segmentation latent space, we propose a two-stage pipeline for decomposable part-level generation, which consists of a whole-level diffusion transformer (DiT) [peebles2023scalable] that generates the global geometry-segmentation latent and a part-level DiT to refine each part latent for complete part geometry of quality.
@@
Our global-level DiT employs a rectified flow model [lipman2022flowmatching], which uses a linear interpolation forward process $Z_{t}=(1-t)Z_{0}+t\epsilon$ to construct the noisy Geom-Seg VecSet latent $Z_{t}$ from the data sample $Z_{0}$ and noise $\epsilon\in\mathbb{R}^{L\times d}$ at a timestep $t\in[0,1]$. In the backward process, our DiT parameterized by $\theta$ learns to model the timestep-dependent velocity field, $v(Z_{t},t)=\delta_{t}Z_{t}$, representing the moving velocity from the noise to the data sample. $\theta$ is optimized on the conditional flow matching (CFM) objective [lipman2022flowmatching]:
@@ 3.3 Part-level Latent Diffusion
Our part-level latent diffusion model adopts a dual-space generation architecture that jointly generates each part’s geometric latent from two spaces. For the $i$-th part, we denote its dual-space latent as $X_{i}^{*}\coloneqq(X_{i}^{\text{gcs}},X_{i}^{\text{ncs}})\in\mathbb{R}^{L\times 2d}$. Specifically, $X_{i}^{\text{gcs}}$ denotes the part latent in the global coordinate space, encoding the part’s relative scale, position, and compositional relationship to the whole object. In contrast, $X_{i}^{\text{ncs}}$ denotes the part latent in a normalized canonical space, capturing high-fidelity part geometry within a normalized $0,1]^{3}$ coordinate frame and sharing the same resolution as the whole-object latent (see illustration in Fig. 1). During dual-space diffusion training, a DiT parameterized by $\omega$ predicts $X_{i}^{*}$ for the $i$-th part, conditioned on the input image $I$, the part latent $X_{i}$ (padded to length $L$), and the whole-object latent $Z_{0}$ (replaced with $\hat{Z_{0}}$ at inference time). The prediction follows the same flow-matching formulation as in the whole-level latent diffusion, so we omit the detailed equations for brevity. To enable the transformer to distinguish tokens from the two geometric contexts, we inject a learnable space embedding $e^{s}\in\mathbb{R}^{1\times d}$, where $s\in\{\text{gcs},\text{ncs}\}$ indexes the coordinate space. The corresponding space embedding is added to $X_{i}^{\text{gcs}}$ and $X_{i}^{\text{ncs}}$ with broadcasting before feeding them into the transformer. Furthermore, we also adopt a global-local attention mechanism inspired by previous work [chen2025ultra3d, yan2025xpart]:

### v2

@@ 1 Introduction
With the rapid development of 3D content creation in fields such as computer graphics, robotics, virtual reality, and digital manufacturing, the demand for controllable and semantically structured 3D shape generation has increased significantly. Traditional 3D generation methods [42, 44, 16, 36] primarily focus on producing holistic meshes that capture only the overall geometry of objects. However, real-world applications—such as part-aware editing, physical simulation, robotic manipulation, and modular design—require not only accurate geometry but also a decomposable understanding of object parts and their relationships. This need has motivated increasing interest in part-level 3D shape generation, which aims to synthesize objects as sets of coherent and semantically meaningful components. Compared to holistic 3D reconstruction, part-level 3D shape generation poses two principal challenges. First, achieving structurally and semantically coherent part segmentation is inherently challenging. The model must infer an appropriate segmentation granularity while preserving both global structural integrity and semantic consistency. The wide variability in object geometries and part compositions across—and even within—categories further makes coherent and meaningful decomposition highly nontrivial. Second, part geometry degradation frequently occurs during generation, particularly for small or thin components. These fine structures tend to lose geometric fidelity due to limited generative resolution and the difficulty of jointly preserving global coherence and local geometric details.
@@
Existing solutions develop implicit segmentation by composing groups of latent vectors for decomposable part generation [18, 31, 4], but suffer from limited controllability over segmentation granularity and suboptimal part-level geometric fidelity. Another line of work first performs whole-object generation and subsequently applies explicit segmentation using multi-view images [3], dense points [37], low-resolution voxels [40], or mesh faces [38]. These approaches achieve higher segmentation quality and controllability but require a powerful auxiliary segmenter trained on abundant labeled data. We begin with the observation that part awareness emerges implicitly during whole-geometry learning. This arises from two factors: (i) intra- and inter-part geometric correlations are implicitly promoted during VAE latent-space construction for efficient geometric compression, leading to part-aware latent representations; (ii) image conditions encoded by pretrained vision models such as DINOv2 [24] carry rich part-level semantic cues, which facilitate part-aware learning in image-to-geometry diffusion. Motivated by these observations, we propose to construct a unified 3D geometry–segmentation latent space based on the VecSet representation [42]. This latent space is learned to jointly encode geometric shape and part-segmentation information into Geom-Seg VecSet, and to decode each latent vector into its corresponding geometric contribution and part label. Thanks to the tight integration of part understanding and geometry learning, our representation maintains reconstruction quality comparable to the original VecSet.
@@
For image-guided part-level 3D generation, we introduce a two-stage latent diffusion model, UniPart. We first perform whole-object diffusion to generate Geom-Seg VecSet, yielding both the global geometric latent and the corresponding part latent mask. Compared with methods that conduct part-level generation with implicit segmentation, our approach produces explicit part-latent masks, providing more flexible segmentation granularity and higher part-level controllability. Unlike previous generate-then-segment pipelines, our approach performs geometry generation and part segmentation jointly. This enables more efficient part learning by leveraging the rich part-aware priors implicitly captured in the geometry generator. Moreover, latent-space segmentation actually strikes a balance between granularity—akin to dense point-based approaches—and robustness—similar to low-resolution voxel methods. In the second stage, we directly use the segmented part latents, along with the global latent, as dual conditioning signals for part-level generation. Compared with prior approaches that require whole-object geometry generation and segmentation on explicit representations to condition the part-level diffusion [37, 38, 40], our latent-based conditioning offers three key advantages: (i) avoiding geometry degradation introduced by decode–encode cycles (e.g., decoding to SDF for lossy marching-cube mesh extraction and re-encoding sampled surface points [37, 38]); (ii) providing domain-aligned conditioning signals that facilitate latent-space diffusion learning; (iii) injecting informative part–whole relational cues inherently preserved in the latent space. Furthermore, our part-level diffusion performs a novel dual-space latent generation procedure that jointly predicts part geometry latents in both the global coordinate space and a normalized canonical space. Decoding these two latents provides the part’s canonical geometry (from the normalized space latent) and its scale and placement within the object (from the global-space latent), which substantially mitigates part-geometry degradation and enables seamless and consistent part composition. Extensive experiments are conducted to validate the effectiveness and superiority of our UniPart on both part segmentation and geometry quality.
@@ 2 Related Work
Recent object-level 3D generation methods fall into two main categories: 2D-lifting and native 3D approaches. DreamFusion [26] pioneered the use of 2D diffusion priors via Score Distillation Sampling [1, 33] to optimize 3D representations, while follow-ups [22, 34, 21, 28] improve multi-view consistency by generating aligned image sets before reconstruction. In parallel, native 3D generative models are trained directly on large-scale datasets [10, 8]. CLAY [42] and related works [44, 16, 15] adopt a VecSet-based VAE–LDM framework [41] to generate meshes and require high-quality watertight meshes for training. Trellis [35] introduces an explicit voxel-based latent space trained with rendering losses, offering stronger geometric inductive bias [17, 13]. However, scaling such representations to high resolution remains a key challenge for fine-detail synthesis. Our work operates at the part level and builds upon the VecSet paradigm: we retain the VecSet-based latent architecture but extend the purely geometric latent space by incorporating geometry-aligned segmentation features, thereby constructing a novel unified geometry–segmentation latent space that enables structured part decomposition while preserving reconstruction fidelity.
@@
Recent part segmentation methods leverage either 2D priors or native 3D supervision. Approaches such as SaMesh [30] lift 2D SAM [14, 27] predictions to 3D via multi-view post-processing, while SAMPart3D [39] and PartField [20] distill SAM priors into a 3D semantic feature field for clustering-based decomposition. In contrast, 3D native methods like P3-SAM [23] and PartSAM [45] are trained directly on large-scale part-annotated meshes and produce point-wise semantic features for mask prompting. GeoSAM2 [11] adopts a hybrid strategy: it finetunes the pretrained SAM on 3D data to enable multi-view consistent segmentation, then back-projects the 2D segmentation onto the mesh surface to obtain a coherent 3D part decomposition. Unlike these strategies that conduct segmentation on explicit 2D/3D representations, our method performs decomposition on the latent representation. By integrating high-dimensional, geometry-aligned semantic features into the latent representation, our approach inherits the robustness and boundary-awareness of sampling while enabling end-to-end part-level generation without relying on hard segmentation masks. Meanwhile, solving the segmentation task as a diffusion-sampling task is much more reasonable than regression. Unlike regression, which forces a single deterministic output, treating part segmentation as a diffusion-based sampling task better captures the inherent ambiguity and multi-modality of valid part decompositions.
@@
Existing part-level 3D generation methods mainly fall into two groups based on the diffusion and autoregressive paradigms. Diffusion-based approaches can be further categorized by their generation strategy: one-shot-for-all methods such as PartCrafter [18], PartPacker [31], OmniPart [40], BANG [43], FullPart [12] and X-Part [37], which jointly generate all parts in a single forward pass; one-shot-per-part models such as HoloPart [38] generates each part independently but conditioned on global context, and AutoPartGen [4] adopts a sequential generation scheme. In contrast, AR-based methods operate directly on mesh sequences: MeshCoder [7] leverages an LLM to produce Blender scripts that are executed to reconstruct parts, and Mesh Silksong [29] employs a manifold-aware mesh representation for autoregressive part synthesis. Except for adopting joint geometry-segmentation diffusion at the first stage, our part-level diffusion is also different in: (i) using latent-based conditions rather than explicit representations, which avoids geometry degradation, provides domain-aligned conditioning signals, and also injects informative part–whole relational cues; (ii) performing a dual-space latent generation procedure that achieves full-resolution part geometry of high quality and enables seamless and consistent part composition.
@@ 3.1 Geometry-Segmentation VAE
[42] is a popular representation that encodes the geometry of a 3D mesh into the latent space. Specifically, dense points $P\in\mathbb{R}^{C\times 6}$ (each of $C$ points is represented by its position and normal vector on the mesh surface) are first sampled from the mesh $\mathcal{O}$. A fixed-length set of latent vectors $Z\in\mathbb{R}^{L\times d}$ is computed through a cross-attention layer between $P$ and its downsampled $L$ points (or a learnable query set) and several self-attention layers, encoded into $d$-dimension latent vectors. This process is conducted by the encoder $\mathcal{E}$ of a transformer-based VAE. The decoder $\mathcal{D}$ then decodes $Z$ into an implicit field (e.g. occupancy field or signed-distance field) for mesh reconstruction, as $\mathcal{D}(Z,q)\rightarrow[0,1]$, where $q$ is the set of query points in the target 3D space. The whole VAE is trained by minimizing the loss function formulated as
@@
Classic VecSet representations encode rich geometric information into the latent space. To better support compositional generation, we further propose Geom-Seg Vecset that additionally encodes part segmentation information into the latent space and conducts latent segmenting in the decoding stage for part-wise representation. In doing so, we first segment object meshes into $N$ parts according to both geometric and semantic patterns. Then $C$ points with part segmentation label are sampled, denoted as $P\in\mathbb{R}^{C\times 7}$, as the input of encoding, derieving the encoded latent $Z=\mathcal{E}(P)\in\mathbb{R}^{L\times d}$. A decoder $\mathcal{D}_{\text{geom}}$ as in VecSet is adopted for geometry reconstruction from $Z$. We further develop a segmentation decoder $\mathcal{D}_{\text{seg}}$ and follow the promptable segmentation pipeline of SAM2 [27], which accepts $Z$ and latent indices $q$ as input and generates segmentation masks that the prompt belongs to. The Geom-Seg VecSet VAE is trained by minimizing the loss function formulated as
@@
where $L_{\text{seg}}$ is the segmentation loss following SAM2 [27]. Considering geometric cues often delineate semantic boundaries, the geometric features and part segmentation patterns of objects are inherently interdependent. It critically facilitates coupled representation learning of geometry and segmentation information. In practice, we found that the Geom-Seg VecSet space can be smoothly learned by fine-tuning a pure geometric VecSet VAE, with the latent space and neural network kept in the original scale (equal in $L$ and $d$), which is sufficient to jointly represent the segmentation and geometry without loss of quality (see visualizations in Sec. 4.3).
@@ 3.2 Whole-level Latent Diffusion
With the learned geometry-segmentation latent space, we propose a two-stage pipeline for decomposable part-level generation, which consists of a whole-level diffusion transformer (DiT) [25] that generates the global geometry-segmentation latent and a part-level DiT to refine each part latent for complete part geometry of quality.
@@
Our global-level DiT employs a rectified flow model [19], which uses a linear interpolation forward process $Z_{t}=(1-t)Z_{0}+t\epsilon$ to construct the noisy Geom-Seg VecSet latent $Z_{t}$ from the data sample $Z_{0}$ and noise $\epsilon\in\mathbb{R}^{L\times d}$ at a timestep $t\in[0,1]$. In the backward process, our DiT parameterized by $\theta$ learns to model the timestep-dependent velocity field, $v(Z_{t},t)=\delta_{t}Z_{t}$, representing the moving velocity from the noise to the data sample. $\theta$ is optimized on the conditional flow matching (CFM) objective [19]:
@@ 3.3 Part-level Latent Diffusion
Our part-level latent diffusion model adopts a dual-space generation architecture that jointly generates each part’s geometric latent from two spaces. For the $i$-th part, we denote its dual-space latent as $X_{i}^{*}\coloneqq(X_{i}^{\text{gcs}},X_{i}^{\text{ncs}})\in\mathbb{R}^{L\times 2d}$. Specifically, $X_{i}^{\text{gcs}}$ denotes the part latent in the global coordinate space, encoding the part’s relative scale, position, and compositional relationship to the whole object. In contrast, $X_{i}^{\text{ncs}}$ denotes the part latent in a normalized canonical space, capturing high-fidelity part geometry within a normalized $0,1]^{3}$ coordinate frame and sharing the same resolution as the whole-object latent (see illustration in Fig. 1). During dual-space diffusion training, a DiT parameterized by $\omega$ predicts $X_{i}^{*}$ for the $i$-th part, conditioned on the input image $I$, the part latent $X_{i}$ (padded to length $L$), and the whole-object latent $Z_{0}$ (replaced with $\hat{Z_{0}}$ at inference time). The prediction follows the same flow-matching formulation as in the whole-level latent diffusion, so we omit the detailed equations for brevity. To enable the transformer to distinguish tokens from the two geometric contexts, we inject a learnable space embedding $e^{s}\in\mathbb{R}^{1\times d}$, where $s\in\{\text{gcs},\text{ncs}\}$ indexes the coordinate space. The corresponding space embedding is added to $X_{i}^{\text{gcs}}$ and $X_{i}^{\text{ncs}}$ with broadcasting before feeding them into the transformer. Furthermore, we also adopt a global-local attention mechanism inspired by previous work [5, 37]: